Calculate the molality of a 6.55 kg sample of a solution of the solute CH²Cl²dissolve in the solvent acetone C³H0O if the sample contains 876 g of methylene chloride

1
Expert's answer
2022-04-06T09:01:01-0400

Moles of methylene chloride = "\dfrac{mass}{MM}=\dfrac{876g}{84.93g/mol}" =10.314mol


Volume of solution = \dfrac{mass of mixture}{density of acetone}

="\dfrac{6550g}{0.784g/cm^3} = 8354.59cm^3"


Molarity og solution;

10.314 mole "\to" 8354.59cm3

? "\to" 1000cm3


="\dfrac{10.314mol x 1000cm^3}{8354.59cm^3} = 1.23M"